OpenStudy (pphalke):

A selective college admits only 8% of its applicants. last year 256 applicants were admitted. how many students applied?

OpenStudy (photon336):

set up a ratio do you know how to do this?

OpenStudy (pphalke):

8/100 256/x

OpenStudy (photon336):

\[\frac{ x }{ 256 } = \frac{ 8 }{ 100 }\]

OpenStudy (pphalke):

I got 20.48 but how would I know how many students applied

OpenStudy (pphalke):

100x=2048 divide by 100 and I got 20.48 is that the answer @Photon336

OpenStudy (leenathan):

i think so

OpenStudy (pphalke):

anyone?

OpenStudy (mathstudent55):

8% are admitted. 256 were admitted. That means 256 is 8% of a much larger number.

OpenStudy (mathstudent55):

8 out of 100 is the same as 256 out of x \(\dfrac{8}{100} = \dfrac{256}{x}\)

OpenStudy (pphalke):

3200 @mathstudent55

OpenStudy (mathstudent55):

Yes. Now let's check. 3200 applied. Only 8% were admitted. 8% of 3200 = 0.08 * 3200 = 256. 256 were admitted just as the problem states.